What is the probability of receiving a 6 or a 9 from a deck of 52 cards and an ace or a king?
suppose to be a percent answer

Found 2 solutions by mom+3, Alan3354:
Answer by mom+3(3)

I come up with 31%

Answer by Alan3354(69443)

What is the probability of receiving a 6 or a 9 from a deck of 52 cards and an ace or a king?
suppose to be a percent answer
-------------------
A 6 or a 9 --> 8/52 = 2/13
An ace or a king is the same = 2/13
-------------
For both, it's the product = 4/169
=~ 2.37%